OpenStudy (anonymous):

A restaurant has one type of lemonade that is 12% sugar and another that is 7% sugar. How many gallons of each type does the restaurant need to make 20 gallons of a lemonade mixture that is 10% sugar? A) 12 gallons of the 12% lemonade and 8 gallons of the 7% lemonade B) 10 gallons of the 12% lemonade and 10 gallons of the 7% lemonade C) 8 gallons of the 12% lemonade and 12 gallons of the 7% lemonade D) 2 gallons of the 12% lemonade and 18 gallons of the 7% lemonade

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

@Luigi0210 can you help me please? (:

OpenStudy (anonymous):

@hartnn I need help once again lol

OpenStudy (luigi0210):

I kind of have to leave soon, but I think hartnn can figure it out c:

OpenStudy (anonymous):

okay!

hartnn (hartnn):

let x be the number of gallons of 12% sugar lemonade y = number of gallons of 7% sugar lemonade since we need 20 gallons of a lemonade mixture that is 10% sugar, we first have x+y =20 got this ?

OpenStudy (anonymous):

so x + y = 20 12 + 7

OpenStudy (anonymous):

nevermind 12 and 7 are the percentages not gallons

hartnn (hartnn):

ok, we need one more equation, 12% of sugar x = 0.12 x 7% of sugar y = 0.07y this should equal, 10% of sugar in 20 gallons mixture which is 0.10*20 so we have 0.12x+0.07y = 0.10*20 see if u get this...

OpenStudy (anonymous):

yes I get it

hartnn (hartnn):

now you have 2 equations, 2 unknowns, can you solve for x and y ? x+y =20 0.12x+0.07y = 0.10*20

OpenStudy (anonymous):

am I supposed to put it in slope intercept form to solve for x and y?

hartnn (hartnn):

you can start by using substitution method from 1st equation y = 20 -x substitute this 'y' in 2nd equation 0.12x +0.07(20-x) = 2 can u get x from this?

OpenStudy (anonymous):

okay.. I got 0.12x + 1.4 = 2

hartnn (hartnn):

what about the 0.07*(-x) term ?

OpenStudy (anonymous):

would I subtract 1.4 - 0.12 to get x?

hartnn (hartnn):

0.12x +0.07(20-x) = 2 0.12x + 0.07 * 20 - 0.07*x = 2 0.12x + 1.4 - 0.07x = 2

hartnn (hartnn):

x (0.12 -0.07) = 2-1.4

hartnn (hartnn):

x(0.05) = 0.6 x= 0.6/ 0.05 = 60/5 = 12

hartnn (hartnn):

go through these steps and tell me if you have any doubt in any step

OpenStudy (anonymous):

so the answer would be A?

hartnn (hartnn):

YES.

OpenStudy (anonymous):

lol im sorry math is my worst subject.. it takes me awhile to get it :/

hartnn (hartnn):

no problem, glad you got it :)

OpenStudy (anonymous):

Thank you! (:

hartnn (hartnn):

welcome ^_^